OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5

OpenStudy (anonymous):

i need this answer

OpenStudy (anonymous):

since the required line is parallel to y=-7x-5..... required slope=-7 let (x1,y1)=(1,-3) y-y1=m*(x-x1)